# ERC-SVD: Error-Controlled SVD for Large Language Model Compression

### v3

@@ 4.3 Ablation Study
We present results on C4 [raffel2020exploring] and seven zero-shot reasoning and understanding tasks, as shown in Table 4. It can be observed that either applying REC or PLC alone has already improved performance. Combining them leads to further improvements, making the performance gap even more significant. Results of other compression ratios are presented in Table 8.
@@ 4.4 Compatibility with Quantization
SVD-based LLM compression methods and quantization are two orthogonal techniques. To demonstrate that our method can be integrated with quantization, we adopt GPTQ [frantar2022gptq] to quantize LLaMA-2-7B compressed by our method and SVD-LLM. As shown in Table 12, our method integrates seamlessly with quantization, achieving superior performance compared to SVD-LLM.
@@ 4.5 Efficiency Results
ERC-SVD not only preserves competitive model performance but also achieves substantial inference speedup on hardware. We evaluate the throughput of compressed models on an NVIDIA A100 GPU, shown in Figure 5. Models compressed by ERC-SVD consistently deliver faster inference than the original model. The speedup grows more pronounced as the batch size grows, indicating that ERC-SVD scales more efficiently under larger workloads. Additionally, Appendix A.6 provides further analysis on how ERC-SVD reduces computational complexity.
@@ 5 Conclusion
In this work, we propose ERC-SVD, a novel post-training SVD-based compression method for LLMs, formulated from an error-controlled perspective. ERC-SVD effectively leverages the residual matrix resulting from SVD truncation to reduce the truncation loss and enhance layer-wise reconstruction accuracy. Furthermore, it selectively compresses only the last few layers of the model, thereby significantly mitigating error propagation. Extensive experiments across various LLM families and benchmark datasets demonstrate that ERC-SVD consistently outperforms existing SVD-based baselines under various settings. These results highlight the effectiveness and generalizability of ERC-SVD in enabling efficient LLM deployment.
@@ A.2 Layer-wise Error Comparison
As discussed in Section 3.2, the same observation also holds here. We conduct experiments on Mistral-7B [jiang2023mistral] and Vicuna-7B [chiang2023vicuna], and the results are presented in Figure 4.
@@ A.3.1 Results on Larger-scale LLMs
In Table 6, we present zero-shot accuracy results for LLaMA-13B [touvron2023llama], LLaMA-30B [touvron2023llama], LLaMA-2-13B [touvron2023llama2], OPT-13B [zhang2022opt], and OPT-30B [zhang2022opt]. Across these evaluation datasets, ERC-SVD consistently outperforms SVD-LLM, with only one minor exception.
@@ A.3.2 Results on Multiple LLM Families
We also evaluate the zero-shot accuracy of OPT-6.7B [zhang2022opt], Mistral-7B [jiang2023mistral], and Vicuna-7B [chiang2023vicuna]. Results are shown in Table 9, and our method consistently outperforms SVD-LLM across these diverse LLM architectures. In addition, we conduct experiments on Qwen-3-8B [yang2025qwen3], with perplexity and zero-shot accuracy results reported in Table 7.
@@ A.3.3 Additional Ablation Results
The ablation results for LLaMA-2-7B [touvron2023llama2] under 40%, 50%, and 60% compression ratios are presented in Table 8. A similar trend can also be observed here: incorporating both REC and PLC leads to a substantial reduction in perplexity across all settings.
Table 8: Ablation results of residual compensation and partial-layer compression on LLaMA-2-7B [touvron2023llama2] under 40%, 50%, and 60% compression ratios.
@@ A.3.4 Impact of calibration data
Table 12: Perplexity ($\downarrow$) of LLaMA-2-7B compressed by ERC-SVD and SVD-LLM, followed by quantization with GPTQ-8bit [frantar2022gptq]. Blue arrows within parentheses highlight the relative improvement.
@@ A.4 Comparison with Smaller-Scale LLM
To further assess the effectiveness of ERC-SVD under relatively high compression ratios, we conduct a comparative evaluation of 50% and 60% compressed LLaMA-2-7B against Gemma-2-2B [gemma_2024]. Table 13 presents the performance comparison between the compressed LLaMA-2-7B models and Gemma-2-2B. Under 50% compression ratio (approximately 3.5B parameters), LLaMA-2-7B consistently outperforms Gemma-2-2B across most benchmarks. Even under a more aggressive 60% compression ratio (around 2.8B parameters), the compressed model maintains competitive performance, remaining comparable to Gemma-2-2B.
@@ A.5 Comparison with Pruning
Table 10 shows the performance of LLaMA-2-7B compressed by LLM-Pruner [ma2023llm] and ERC-SVD under different compression ratios on the WikiText-2 dataset. It can be observed that ERC-SVD consistently outperforms the pruning method, achieving a perplexity of 58.88 under 60% compression ratio, compared to 114.23 for LLM-Pruner.
@@ A.7 Demonstration of Generated Contents
Table 15: Generated contents of LLaMA-7B [touvron2023llama] compressed by ERC-SVD.
@@ A.8 ERC-SVD on Vision-Language Model
In this section, we apply ERC-SVD to the language part of vision-language models (VLMs) such as LLaVA [liu2024improved], and evaluate both quantitative and qualitative performance.
@@ A.8.1 Quantitative Results
We compare the performance of LLaVA-1.5-7B [liu2024improved] compressed using SVD-LLM [wang2024svd] and ERC-SVD under 20% compression ratio. We report results on several benchmarks: POPE [li2023evaluating], TextVQA [singh2019towards], and ScienceQA [lu2022learn]. The results, shown in Table 14, indicate that ERC-SVD consistently outperforms SVD-LLM across all benchmarks. Notably, it achieves substantial relative improvements, 66% on TextVQA and 40% on ScienceQA. Moreover, on the POPE-random and POPE-popular subsets, the model compressed by ERC-SVD even surpasses the original LLaVA-1.5-7B.

### v4

@@ 4.3 Ablation Study
We present results on C4 [43] and seven zero-shot reasoning and understanding tasks, as shown in Table 4. It can be observed that either applying REC or PLC alone has already improved performance. Combining them leads to further improvements, making the performance gap even more significant. Results of other compression ratios are presented in Table 8.
@@ 4.4 Compatibility with Quantization
SVD-based LLM compression methods and quantization are two orthogonal techniques. To demonstrate that our method can be integrated with quantization, we adopt GPTQ [12] to quantize LLaMA-2-7B compressed by our method and SVD-LLM. As shown in Table 12, our method integrates seamlessly with quantization, achieving superior performance compared to SVD-LLM.
@@ 4.5 Efficiency Results
ERC-SVD not only preserves competitive model performance but also achieves substantial inference speedup on hardware. We evaluate the throughput of compressed models on an NVIDIA A100 GPU, shown in Figure 5. Models compressed by ERC-SVD consistently deliver faster inference than the original model. The speedup grows more pronounced as the batch size grows, indicating that ERC-SVD scales more efficiently under larger workloads.
@@ 5 Conclusion
In this work, we propose ERC-SVD, a novel post-training SVD-based compression method for LLMs, formulated from an error-controlled perspective. ERC-SVD effectively leverages the residual matrix resulting from SVD truncation to reduce the truncation loss and enhance layer-wise reconstruction accuracy. Furthermore, it selectively compresses only the last few layers of the model, thereby significantly mitigating error propagation. Extensive experiments across various LLM families and benchmark datasets demonstrate that ERC-SVD consistently outperforms existing SVD-based baselines under various settings.
Acknowledgement
This paper is supported by Young Scientists Fund of the National Natural Science Foundation of China (NSFC) (No. 62506305), Zhejiang Leading Innovative and Entrepreneur Team Introduction Program (No. 2024R01007), Key Research and Development Program of Zhejiang Province (No. 2025C01026), Scientific Research Project of Westlake University (No. WU2025WF003), Chinese Association for Artificial Intelligence (CAAI) & Ant Group Research Fund - AGI Track (No. 2025CAAI-ANT-13). It is also supported by the research funds of the National Talent Program and Hangzhou Municipal Talent Program.
@@ A.2 Layer-wise Error Comparison
As discussed in Section 3.2, the same observation also holds here. We conduct experiments on Mistral-7B [39] and Vicuna-7B [40], and the results are presented in Figure 4.
@@ A.3.1 Results on Larger-scale LLMs
In Table 6, we present zero-shot accuracy results for LLaMA-13B [4], LLaMA-30B [4], LLaMA-2-13B [37], OPT-13B [38], and OPT-30B [38]. Across these evaluation datasets, ERC-SVD consistently outperforms SVD-LLM, with only one minor exception.
@@ A.3.2 Results on Multiple LLM Families
We also evaluate the zero-shot accuracy of OPT-6.7B [38], Mistral-7B [39], and Vicuna-7B [40]. Results are shown in Table 9, and our method consistently outperforms SVD-LLM across these diverse LLM architectures. In addition, we conduct experiments on Qwen-3-8B [6], with perplexity and zero-shot accuracy results reported in Table 7.
@@ A.3.3 Additional Ablation Results
The ablation results for LLaMA-2-7B [37] under 40%, 50%, and 60% compression ratios are presented in Table 8. A similar trend can also be observed here: incorporating both REC and PLC leads to a substantial reduction in perplexity across all settings.
Table 8: Ablation results of residual compensation and partial-layer compression on LLaMA-2-7B [37] under 40%, 50%, and 60% compression ratios.
@@ A.3.4 Impact of calibration data
Table 12: Perplexity ($\downarrow$) of LLaMA-2-7B compressed by ERC-SVD and SVD-LLM, followed by quantization with GPTQ-8bit [12]. Blue arrows within parentheses highlight the relative improvement.
@@ A.4 Comparison with Smaller-Scale LLM
To further assess the effectiveness of ERC-SVD under relatively high compression ratios, we conduct a comparative evaluation of 50% and 60% compressed LLaMA-2-7B against Gemma-2-2B [51]. Table 13 presents the performance comparison between the compressed LLaMA-2-7B models and Gemma-2-2B. Under 50% compression ratio (approximately 3.5B parameters), LLaMA-2-7B consistently outperforms Gemma-2-2B across most benchmarks. Even under a more aggressive 60% compression ratio (around 2.8B parameters), the compressed model maintains competitive performance, remaining comparable to Gemma-2-2B.
@@ A.5 Comparison with Pruning
Table 10 shows the performance of LLaMA-2-7B compressed by LLM-Pruner [16] and ERC-SVD under different compression ratios on the WikiText-2 dataset. It can be observed that ERC-SVD consistently outperforms the pruning method, achieving a perplexity of 58.88 under 60% compression ratio, compared to 114.23 for LLM-Pruner.
@@ A.7 Demonstration of Generated Contents
Table 15: Generated contents of LLaMA-7B [4] compressed by ERC-SVD.
@@ A.8 ERC-SVD on Vision-Language Model
In this section, we apply ERC-SVD to the language part of vision-language models (VLMs) such as LLaVA [50], and evaluate both quantitative and qualitative performance.
@@ A.8.1 Quantitative Results
We compare the performance of LLaVA-1.5-7B [50] compressed using SVD-LLM [28] and ERC-SVD under 20% compression ratio. We report results on several benchmarks: POPE [52], TextVQA [53], and ScienceQA [54]. The results, shown in Table 14, indicate that ERC-SVD consistently outperforms SVD-LLM across all benchmarks. Notably, it achieves substantial relative improvements, 66% on TextVQA and 40% on ScienceQA. Moreover, on the POPE-random and POPE-popular subsets, the model compressed by ERC-SVD even surpasses the original LLaVA-1.5-7B.
@@ A.8.2 Qualitative Results
A.9 Future Work
While this work focuses on general LLM compression, the low-rank structures identified via SVD hold significant potential for reasoning LLMs [55, 6, 56]. Beyond autoregressive language tasks, the principles of SVD-based compression are readily transferable to diffusion frameworks [57, 58, 59]. In related tasks, the Transformer-based denoisers often exhibit extreme parameter redundancy. By leveraging ERC-SVD to decompose attention layers, maybe we can accelerate the iterative sampling process, making real-time deployment of generative models on edge devices more feasible without sacrificing structural fidelity. Furthermore, the intersection of model compression and hallucination mitigation [60, 61] presents an intriguing research frontier. We hypothesize that low-rank approximations could potentially filter out the noise that contributes to factually incorrect generations, a direction we intend to explore in future work.